# Quantum Monte Carlo description of correlated electrons in two-dimensional FeSe

### v1

@@ III Numerical Results
Table 2: VMC and DMC energies of double layer FeSe obtained using determinant single-particle orbital basis set $8s6pnd1f;(n=4,6)$ and $6s4p2d$ for $Fe$ and $Se$, respectively, and Jastrow orbital $3s3pnd,(n=1,3)$ and $3s2p$ for Fe and Se, respectively. On the number of $d$ orbital is listed in the first column. Energies are in Ha/el and are calculated using two-body Jastrow $u_{F}$. The energies are obtained for double-layer FeSe with Fe-Fe bond length 3.689 Åand the distance between layers 5.83 Å.
Figure 3: Energy of double-layer FeSe with respect to the minimum energy $E_{min}$ as a function of separation between layers. The energies are calculated using VMC and DMC with JDFT and JSD wave functions and also DFT. The energy curves are obtained for double-layer FeSe with Fe-Fe bond length 3.689Å(top panel) and 4.233Å(bottom panel).
We calculated the VMC and DMC ground state energy curve of double-layer FeSe as a function of interlayer separation. We considered the experimental inplane iron-iron bond length of 3.689 Åand the stretched iron-iron bond length of 4.233Å. The energy curves, which are obtained using the $JDFT$ and $JSD$ wave functions, are illustrated in Fig. 3. The DFT, VMC, and DMC energy values of the data points in Fig. 3, are listed in tables. 3 and 4. All energy data points in Fig. 3 are obtained using the basis set $8s6p4d1f/3s3p1d$ for the iron atom.
$d$(Å) $HF$ $DFT$ $VMC-JDFT$ $VMC-JSD$ $DMC-JDFT$ $DMC-JSD$
$4.27$ $-5.3439236$ $-6.0112733$ $-6.021038(6)$ $-6.027753(5)$ $-6.033048(9)$ $-6.035778(8)$
$4.79$ $-5.3472775$ $-6.0118530$ $-6.021618(6)$ $-6.028438(5)$ $-6.033569(9)$ $-6.036546(8)$
$5.16$ $-5.3499789$ $\mathbf{-6.0121258}$ $-6.022014(6)$ $-6.028825(5)$ $\mathbf{-6.034169(9)}$ $-6.037144(8)$
$5.53$ $-5.3518114$ $-6.0118895$ $-6.021896(6)$ $-6.028921(5)$ $\mathbf{-6.034165(9)}$ $-6.037319(8)$
$5.83$ $-5.3531499$ $-6.0114938$ $\mathbf{-6.022123(6)}$ $\mathbf{-6.029479(5)}$ $-6.033987(9)$ $\mathbf{-6.037408(8)}$
$6.09$ $-5.3538347$ $-6.0110016$ $-6.020785(6)$ $-6.029394(5)$ $-6.033591(9)$ $-6.036983(8)$
$6.27$ $-5.3542681$ $-6.0105423$ $-6.020053(6)$ $-6.028906(5)$ $-6.033299(9)$ $-6.036674(8)$
Table 3: Ground state energy of double layer FeSe as a function of interlayer separation ($d$) in Ha/el. The energies are obtained using a simulation cell with $176$ electrons, two-body Jastrow term $J_{F}$, and Fe-Fe bond-length 3.651 Å.
$d$(Å) $HF$ $DFT$ $VMC-JDFT$ $VMC-JSD$ $DMC-JDFT$ $DMC-JSD$
$3.39$ $-5.3454731$ $-6.008139$ $-6.018283(6)$ $-6.025481(6)$ $-6.03039(5)$ $-6.03332(5)$
$3.81$ $-5.3497955$ $\mathbf{-6.009011}$ $-6.019812(6)$ $-6.027055(5)$ $-6.03180(5)$ $-6.03487(4)$
$4.23$ $-5.3522866$ $-6.008861$ $\mathbf{-6.020030(6)}$ $\mathbf{-6.027989(5)}$ $\mathbf{-6.03202(5)}$ $\mathbf{-6.03571(4)}$
$4.66$ $-5.3539755$ $-6.008240$ $-6.019513(6)$ $-6.027859(5)$ $-6.03177(5)$ $\mathbf{-6.03574(4)}$
$5.08$ $-5.3550693$ $-6.007422$ $-6.018945(6)$ $-6.027605(5)$ $-6.03137(5)$ $-6.03556(5)$
$5.50$ $-5.3557858$ $-6.006575$ $-6.018307(6)$ $-6.027461(5)$ $-6.03089(5)$ $-6.03552(5)$
$5.93$ $-5.3561037$ $-6.005549$ $-6.017355(6)$ $-6.026636(5)$ $-6.03033(5)$ $-6.03485(5)$
$6.35$ $-5.3559713$ $-6.004349$ $-6.016710(6)$ $-6.026427(5)$ $-6.02962(5)$ $-6.03455(5)$
$6.69$ $-5.3555844$ $-6.003223$ $-6.015909(6)$ $-6.025875(5)$ $-6.02874(5)$ $-6.03386(4)$
$6.98$ $-5.3552118$ $-6.002385$ $-6.014770(6)$ $-6.024265(5)$ $-6.02820(5)$ $-6.03273(5)$
Table 4: Ground state energy of double layer FeSe as a function of interlayer separation ($d$) in Ha/el. The energies are obtained using a simulation cell with $176$ electrons, two-body Jastrow term $J_{F}$, and Fe-Fe bond-length 4.233 Å.
Figure 4: (Top panel) Correlation energy per electron for double layer FeSe as a function of interlayer separation. The energies are obtained using VMC and DMC with $JSD$ wave function for two Fe-Fe bond-length ($bl$) 3.689 and 4.233 Å. (Bottom panel) Geometry of double layer FeSe in which d is the distance between the center of mass of each layer. Fe and Se atoms are brown and green spheres, respectively.
The VMC and DMC correlation energy of the double layer FeSe as a function of the distance between the layers is shown in Fig. 4. The results were calculated using the wave function $JSD$ which is more accurate than $JDFT$. The correlation energies were obtained for two different Fe-Fe bond lengths 3.689 and 4.233 Åand system size 176 electrons.
IV Discussion
We specifically focused on optimizing the trial wave function for each system size and lattice parameter before performing the final VMC and DMC simulations. The trial wave function is crucial in QMC simulations because it directly affects the accuracy, efficiency, and stability of the simulations, as it serves as a guiding function for importance sampling in DMC, which modifies the diffusion equation to improve convergence. A properly optimized trial wave function can lead to large statistical fluctuations, making it difficult to obtain precise results. Optimizing the trial wave function reduces variance, leading to faster and more stable simulations. We systematically optimized the trial wave function by following three steps: (I) The Jastrow coefficients and the two-body term $u$ were optimized, as shown in Fig. 1(a). In this step, depending on the size of the system, $\sim 4-8$ mHa / el VMC energy obtained with respect to DFT. (II) The Jastrow function was fully optimized (Fig. 1(b)) and the wave function was called $JDFT$. The energy difference between VMC and DFT is $\sim 8-11$ mHa/el, depending on the system size. (III) The Jastrow and Slater determinant were optimized simultaneously and the wave function was named $JSD$. Since the FS errors normally increase by the number of electrons in the simulation cell, the VMC energy gain with respect to DFT is smaller for larger systems. We found that following these three steps in order, meaning that the well-optimized wave function at the end of each step is used as the initial wave function for the next step, is required to reach an accuracy of mHa/el and avoid local minima.
Performing QMC simulations in a finite supercell with periodic boundary conditions (PBC) suffers from finite-size (FS) errors [54]. The main source of FS errors includes (i) the single-particle FS effect, which is because of the discrete $k$-point sampling in a finite simulation cell, the electron momentum states are quantized leading to errors in kinetic and exchange energies. (ii) Coulomb FS error, which is due to the issue that the long-range Coulomb interaction is artificially truncated by PBCs and therefore self-interaction and electrostatic image effects distort the interaction energy. (iii) Many-body correlation FS error mainly driven by the absence of long-wavelength collective excitations in small simulation cells. To correct FS errors, we used the Ewald summation and FS scaling extrapolation as shown in Fig. 2. The single-particle FS errors in the kinetic energy part of Hamiltonian can be corrected by twist averaging [54]. The almost perfect linear behavior of the VMC and DMC energies as a function of $1/N$ suggests that the single-particle FS errors are relatively small in our simulations.
@@ IV Discussion
Our VMC and DMC energies for double-layer FeSe at equilibrium configuration are obtained using two forms for two-body Jastrow functions. The pair correlation function $u_{F}(r)$ is more suitable in atomic chemical bond separation, while $u_{C}(r)$, which is large and positive for $r=0$ and goes to zero as $r\rightarrow\infty$, is useful when the atoms are at a distance greater than the variational parameter $b$. In principle, the VMC energies obtained from the two-body Jastrow $u_{F}(r)$ and $u_{C}(r)$ should agree with statistical noise, as our results show (Fig. 2). The VMC energies obtained by the wave function $JSD$ and $u_{F}(r)$ and $u_{C}(r)$ agree completely with the error bars. We propose the idea that using two different forms for the two-body Jastrow term and finding a good agreement between the VMC energies can be used as a test for the accuracy of wave function optimization. The DMC energies are almost independent of the form of Jastrow functions, as our results show (Fig. 2), since the nodal surface is not directly affected by the Jastrow function. The DMC correlation energy of double layer FeSe at equilibrium obtained using the $JDFT$ and $JSD$ wave functions is $0.6789(2)$ and $0.6817(3)$ Ha/el, respectively. The correlation energy of moderately correlated transition metals such as Fe and Ni is $\sim 0.2-0.4$ Ha/el [55], while the strongly correlated transition metal oxides such as FeO and NiO have the correlation energy of $\sim 0.5-1.0$ Ha/el [40, 56]. Our DMC results indicate that the el-el correlation energy of double-layer FeSe is in the strongly correlated range. The bulk FeSe with the tetragonal crystal structure is metallic [57]. Hence, it would be expected that electrons are moderately correlated. Our DMC results predict that the three- to two-dimensional transition increases the el-el correlation energy in FeSe. Therefore, the localised Hubbard $U$ parameter, widely used in DFT+$U$ simulations for transition metal compounds, for double-layer FeSe is larger than the one for crystal FeSe.
Comparison of the energy curves as a function of the distance between FeSe layers (Fig. 3) shows that the optimized separation between FeSe layers decreases as the iron-iron bond length stretches. The energy curves $VMC$-$JDFT$ and $DMC$-$JDFT$ for Fe-Fe = 3.689 Åshow a degeneracy near the minimum which can be relatively removed by optimizing the Slater determinant as the energy curves $VMC$-$JSD$ and $DMC$-$JSD$ display. However, this degeneracy vicinity of the energy minimum becomes more pronounced with increasing iron-iron bond length, and cannot be removed by optimizing the Slater determinant as shown by the energy curves $VMC$-$JSD$ and $DMC$-$JSD$ for Fe-Fe = 4.233Å. These results indicate that stretching the bond length in the strongly correlated double-layer FeSe system enlarges the static correlation driven by the localization of electrons of the iron d-orbital. Hence, using multideterminant or at least using better single-particle orbitals than DFT-LDA such as multiconfigurational self-consistent field orbitals may improve the description of the static correlation effects in this system.
The VMC and DMC results with the wave function $JSD$ show that the absolute value of the correlation energy decreases as the distance between the layers of FeSe double-layer increases (Fig. 4). Within the interlayer separation range studied, the absolute value of the correlation energy of the FeSe double layer with a Fe-Fe bond length 3.689 Åat fixed $d$ is higher than in the system with Fe-Fe bond length 4.233 Å. However, the correlation energy of the system with the Fe-Fe bond length 3.689 Ådecreases faster with $d$ than the stretched system. The extrapolation of our results predicts that the correlation energy of two systems at $d>7$may cross over. This suggests that in-plane stretch affects the nature of interaction between layers in double-layer FeSe.
Acknowledgements

### v2

@@ III.2 Correlation energy of Double-layer FeSe at the thermodynamic limit
Table 5: VMC and DMC energies of double layer FeSe obtained using determinant single-particle orbital basis set $8s6pnd1f;(n=4,6)$ and $6s4p2d$ for Fe and Se, respectively, and Jastrow orbital $3s3pnd,(n=1,3)$ and $3s2p$ for Fe and Se, respectively. On the number of $d$ orbital is listed in the first column. Energies are in Ha/el and are calculated using two-body Jastrow $u_{F}$. The energies are obtained for double-layer FeSe with Fe-Fe bond length 3.69 Å and the distance between layers 5.83 Å. The number of electrons in the simulation cell is 176.
@@
Refer to caption Refer to caption
Figure 4: Energy of double-layer FeSe with respect to the minimum energy $E_{min}$ as a function of separation between layers $d$. The energies are calculated using VMC and DMC with JDFT and JSD wave functions and also DFT (LDA). The energy curves are obtained for double-layer FeSe with Fe-Fe bond length (a) 3.69, (b) 4.23, and (c) 4.73 Å.
We compared the correlation energy of double-layer FeSe at the thermodynamic limit with that of isolated Fe and Se atoms, as summarized in Table 6. The correlation energies were computed using ccECP and the $u_{F}$ form for the two-body Jastrow factor. Both VMC and DMC results obtained with the JSD wave function indicate that the average of the atomic correlation energies of Fe and Se provides a reasonable approximation for the correlation energy of the extended FeSe system. This suggests that the total correlation energy of FeSe is primarily determined by the atomic contributions, with the bonding between atoms playing a comparatively minor role in $E_{corr}$. The results in Table 6 show that the magnitude of $E_{corr}$/el of FeSe lies between those of the isolated Fe and Se atoms, being smaller than in Fe but larger than in Se. This trend reflects the nature of the valence orbitals. In fact, Fe $d$ electrons are spatially localized, leading to strong onsite Coulomb interactions, while the $s$- and $p$-valence electrons of Se are more delocalized. A comparison of DMC-JDFT and DMC-JSD calculations for FeSe and Fe further reveals that the difference between DMC-JSD(FeSe) and DMC-JSD(Fe) is more than twice that between DMC-JDFT(FeSe) and DMC-JDFT(Fe). An analogous trend is observed in the VMC results obtained with the JDFT and JSD WFs.
system VMC-JDFT VMC-JSD DMC-JDFT DMC-JSD
FeSe -45.74(1) -52.51(1) -58.50(1) -61.0(2)
Fe -50.02(4) -66.14(3) -64.92(6) -74.72(6)
Se -14.34(3) -44.78(1) -15.07(6) -44.76(3)
Table 6: Correlation energy of bilayer FeSe at the infinite system size limit, Fe and Se atoms in mHa/el obtained using VMC and DMC methods with JDFT and JSD WFs.
III.3 Effects of tensile strain and interlayer separation on correlation energy
In this section, the effects of tensile stress on the geometrical and electronic structure properties of double-layer FeSe are studied. We also calculated the VMC and DMC energy curves of double-layer FeSe as functions of interlayer separation. We considered the experimental inplane iron-iron bond length of 3.69 Å and the stretched iron-iron bond lengths of 4.23, and 4.73 Å. The energy curves, which are obtained using the JDFT and JSD wave functions, are illustrated in Fig. 4. All energy data points in Fig. 4 are obtained using the basis set $8s6p4d1f/3s3p1d$ for the iron atom and N=176 electrons in the simulation cell. Comparison of the energy curves as a function of the distance between FeSe layers (Fig. 4) shows that the optimized separation between FeSe layers decreases as the iron-iron bond length stretches. This suggests that increasing the electron density localization as a result of in-plane stretch enlarges the attractive van der Wall interaction between two layers of double-layer FeSe.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Correlation energy per electron for double layer FeSe as a function of interlayer separation obtained using (a) VMC-JDFT, (b) VMC-JSD, (c) DMC-JDFT, and (d) DMC-JSD for three Fe-Fe bond-length of 3.69, 4.23, and 4.73 Å.
Figure 5 presents the VMC and DMC correlation energies of double-layer FeSe as a function of the interlayer distance. The calculations were performed using JDFT and JSD trial wave functions in a simulation cell containing $N=176$ electrons. Both VMC and DMC results consistently show that the magnitude of the correlation energy decreases with increasing interlayer separation. Moreover, JDFT-based wave functions predict that, for a fixed interlayer spacing $d$, the absolute value of $E_{corr}$ decreases with in-plane stretching. The JSD-based VMC and DMC calculations reveal a similar trend, though with smaller variations in $E_{corr}$ across the studied Fe–Fe distances.
Defining the parameter $\Delta_{corr}=(E_{corr}^{Fe}+E_{corr}^{Se})/2-E_{corr}^{FeSe}$, we find that $\Delta_{corr}$ decreases with increasing interlayer separation as well as with in-plane stretching. This behavior reflects the fact that, upon dissociation of double-layer FeSe into its atomic fragments, $\Delta_{corr}\rightarrow 0$. The rate of change in $\Delta_{corr}$ depends on the wave function and also on the QMC method.
IV Conclusion
We computed the VMC and DMC energies of isolated Fe and Se atoms, as well as of double-layer FeSe, using both JDFT and JSD wave functions. The $E_{\text{corr}}$ obtained for the atoms are in good agreement with the all-electron QC results. For double-layer FeSe, we performed calculations at three different system sizes and extrapolated the energies to the thermodynamic limit to estimate $E_{\text{corr}}$ at infinite system size. Our results indicate that the atomic contributions dominate the correlation energy of double-layer FeSe, while the contribution from bonding is comparatively minor.
We calculated $E_{corr}$ of double-layer FeSe as a function of in-plane tensile strain and interlayer separation. The results reveal a clear trend that the magnitude of $E_{corr}$ decreases with increasing in-plane stretch and interlayer distance, approaching the values corresponding to isolated atomic fragments [60].
Our QMC results indicate that in FeSe and potentially in other transition metal compounds, correlation effects remain large and are not reduced relative to the atomic case. In contrast, for simple metals, where the electron liquid model provides a good description, electron delocalization and screening of the Coulomb interaction generally suppress the correlation energy compared to their atomic counterparts. However, in FeSe and related iron-based systems, the Fe-$3d$ states are spatially localized and exhibit narrow bandwidths relative to the $s$ or $p$ states in simple metals. This localization enhances the el–el Coulomb interactions and consequently increases the correlation energy per electron. An obvious conclusion is that in transition metal compounds, the bands are narrower, so the on-site Coulomb repulsion $U$ is more significant relative to the band-width $W$ making the effective correlation strength $U/W$ larger than in simple metals.
Acknowledgments